# From Rotations to Unitaries: Reversible Quantum Processes and the Emergence of the SU(2)SO(3)SU(2)-SO(3) Isomorphism

### v1

@@ Abstract
We present an operational reconstruction of the well-known (quasi-)isomorphism between the groups $SU(2)$ and $SO(3)$, grounded in the physical description of quantum state preparation and evolution. Starting from the connection between vectors in three-dimensional physical space and quantum states of two-level systems, we investigate how reversible transformations—modeled as CPTP maps—give rise to a correspondence between spatial rotations and unitary operations. Our approach reveals how this group-theoretic structure naturally emerges from physical constraints, particularly the preservation of purity and reversibility in quantum processes. Beyond its theoretical relevance, the construction offers a pedagogically accessible framework for introducing core ideas in quantum mechanics and symmetry groups, making the abstract correspondence between $SU(2)$ and $SO(3)$ tangible through experimentally meaningful procedures.
Keywords: Quantum channels, Group isomorphism, Operational approach.
@@ I Introduction
Physical processes in quantum theory are described by completely positive and trace-preserving (CPTP) maps. By processes, we intend this term to be as broad as possible. Examples include measurement processes via quantum instruments [1], (ir)reversible time evolution [2], coarse-grained descriptions of blurred or saturated detectors [3], LOCC protocols composed of CPTP maps for applications such as entanglement theory [4], and the modeling of open quantum systems’ dynamics, where CPTP maps are employed to reproduce non-unitary evolution due to decoherence and possible dissipation [5]. This list can be extended indefinitely, reflecting the ubiquitous role of such maps in quantum theory and beyond (GTPs) [6].
@@
Moreover, since, once again, we aim to map states to states, it is important to consider that, due to entanglement, local operations on subsystems do not always preserve the positivity of the overall state’s eigenvalues. A standard example is the transpose operation, which is positive but not completely positive [7]. For this reason, we also require the completely positive (CP) part.
Together, these two conditions—which define what is known as a quantum channel (or physical map)—appear to impose strong restrictions on the invertibility of CPTP maps with inverses that are still physically admissible. One might go so far as to say that it is rare, or even contingent, to find a CPTP map whose inverse is also CPTP. For instance, the depolarizing channel has an inverse that is not positive [8].
This question has already been studied in detail [8], where the characterization of invertible quantum channels is addressed via their Kraus representation. In essence, a physical map admits a CPTP inverse if and only if the Kraus decomposition of the former contains exactly one Kraus operator.
In our contribution, we take this analysis a step further by exploring the connection between invertible CPTP maps with physical admissible inverse and group theory. Remarkably, the well-known (quasi-)isomorphism between $SU(2)$ and $SO(3)$ can be recovered in a surprisingly simple way through a preparation procedure involving two-level quantum systems embedded in three-dimensional physical space.
@@
seriously to conclude that any two-level quantum system could, in principle, be prepared in a three-dimensional physical space.
Building on this, we go further to investigate transformations between vectors in $\mathds{D}$ (which give rise to quantum states) and the corresponding allowed state transformations in the space of states. The sequence of these transformations is expected to make the diagram in Fig. 1 commute.
@@
On the one hand, an allowed transformation of a vector in physical space is represented by a rotation, which in turn induces a unitary transformation of the corresponding quantum state. This gives rise to the direction $SO(3)\rightarrow SU(2)$ in the isomorphism we aim to recover. It corresponds to the path represented by the black arrows in the diagram shown in Fig. 1. On the other hand, the reverse direction, $SU(2)\rightarrow SO(3)$, is obtained by requiring a reversible physical transformation of a quantum state. As previously discussed, the only permissible process of this kind is a unitary operation, which translates into a rotation in physical space. This step is illustrated by the blue arrows in the diagram.
To construct both directions in the diagram, we proceed as follows. In the next section, we show how a rotation in physical space unambiguously induces a unitary transformation on the quantum state associated with the corresponding vector. Section III provides a brief review of CPTP maps, with a focus on their Kraus representation. In particular, we discuss a characterization theorem for invertible CPTP maps with physically admissible inverses. Building on this result, in Section IV, we start from a unitary representing a reversible quantum process and reconstruct the associated rotation, thereby recovering the group isomorphism we set out to rediscover. Finally, Section V is left for conclusions and some perspectives.
@@ II From physical spaces to state spaces: $SO(3)\rightarrow SU(2)$
Given a unit vector (called the Bloch vector)
$\hat{r}=(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta),$ (1)
with $\theta\in[0,\pi]$ and $\varphi\in[0,2\pi)$ in the physical space $\mathds{D}\cong\mathds{R}^{3}$, one constructs the pure state
@@
belonging to the so-called Bloch sphere [7]. The parameters $\theta$ and $\varphi$ may have a concrete physical interpretation. For example, in [11] it is shown that they indicate the direction of a magnetic field in the Stern-Gerlach experiment, preparing a bean of spin-up particles in that direction.
@@
With this result in hand, we turn to transformations in physical space that preserve the nature of the quantum state—whether pure or mixed. Such transformations must be isometries, and can be described as rotations around a given axis $\hat{n}$ by an angle $\alpha$, denoted $\mathcal{R}(\hat{n},\alpha)$. Boosts and translations also keep the norm of vectors inert. However both of them do not change the vector as the former is, actually, a representative of an entire class of parallel-displaced ordered pair of points [12]. Although excluded here, such transformations play a central role in exploring the connection of symmetries in quantum mechanics, observables as their generators and group theory, see, for example, [13].
@@
Given a general rotation element $\mathcal{R}(\hat{n},\alpha)\in SO(3)$. The state $\rho^{\prime}$, described by the vector $\vec{r}\,^{\prime}=\mathcal{R}(\hat{n},\alpha)\vec{r}$, can be calculated by applying a matrix $\mathcal{U}(\hat{n},\alpha)$ $\in SU(2)$ on the state $\rho$, described by the original vector $\vec{r}$, in sucha a way that $\rho^{\prime}=\mathcal{U}(\hat{n},\alpha)\rho\,\mathcal{U}^{*}(\hat{n},\alpha)$.
@@
$\mathcal{R}_{jk}(\hat{n},\alpha)$ $=\delta_{jk}\cos\alpha+n_{j}n_{k}(1-\cos\alpha)-\sum^{3}_{l=1}\varepsilon_{jkl}n_{l}\sin\alpha$ (4)
@@
and the group generators $\vec{\tau}$ are given by $(\tau_{l})_{jk}=-i\varepsilon_{jkl}$ [16]. Form now on, Latin letters ($i,j,k,$...), as costumary, assume the values $1,2,3$.
Now, if $\vec{r}\,^{\prime}=\mathcal{R}(\hat{n},\alpha)\vec{r}$, then the rotated vector $\vec{r}\,^{\prime}$ gives rise to a state $\rho^{\prime}$ that we seek to determine. We have
$\rho^{\prime}$ $=\frac{1}{2}\left(\mathds{I}+\vec{r}\,^{\prime}\cdot\vec{\sigma}\right)=\frac{1}{2}\left(\mathds{I}+(\mathcal{R}\vec{r})\cdot\vec{\sigma}\right)$ (6)
$=\frac{1}{2}[\mathds{I}+\cos\alpha\vec{r}\cdot\vec{\sigma}+(1-\cos\alpha)(\hat{n}\cdot\vec{r})(\hat{n}\cdot\vec{\sigma})$ (7)
$-\sin\alpha(\vec{r}\times\hat{n})\cdot\vec{\sigma}]$ (8)
Rearranging terms, with usual trigonometric identities, one finds
$\rho^{\prime}=\mathcal{U}(\hat{n},\alpha)\rho\,\mathcal{U}^{*}(\hat{n},\alpha)$ (9)
@@
$\mathcal{U}(\hat{n},\alpha)=\cos\frac{\alpha}{2}\mathds{I}-i\sin\frac{\alpha}{2}\hat{n}\cdot\vec{\sigma}=e^{-i\frac{\alpha}{2}\hat{n}\cdot\vec{\sigma}},$ (10)
@@
The result above shows that each rotation $\vec{r}\,^{\prime}=\mathcal{R}(\hat{n},\alpha)\vec{r}$, with $\mathcal{R}(\hat{n},\alpha)=e^{-i\alpha\hat{n}\cdot\vec{\tau}}$ induces a state transformation $\rho^{\prime}=\mathcal{U}(\hat{n},\alpha)\rho\,\mathcal{U}^{*}(\hat{n},\alpha)$ with $\mathcal{U}(\hat{n},\alpha)=e^{-i\frac{\alpha}{2}\hat{n}\cdot\vec{\sigma}}$. Thus we have just established the mapping $\phi:SO(3)\rightarrow SU(2)$, with $\phi(\mathcal{R}(\hat{n},\alpha))=\mathcal{U}(\hat{n},\alpha)$. Our next step will be focused on constructing the inverse of $\phi$. To that end, we will make a short digression on CPTP maps.

### v2

# From Rotations to Unitaries: Reversible Quantum Processes and the Emergence of the SU(2)SO(3)SU(2)-SO(3) Homomorphism

@@ Abstract
We present an operational reconstruction of the well-known two-to-one homomorphism between the groups $SU(2)$ and $SO(3)$, grounded in the physical description of quantum state preparation and evolution. Starting from the connection between vectors in three-dimensional physical space and quantum states of two-level systems, we investigate how reversible transformations—modeled as completely positive and trace-preserving maps—give rise to a correspondence between spatial rotations and unitary operations. Our approach reveals how this group-theoretic structure naturally emerges from physical constraints, particularly the preservation of purity and reversibility in quantum processes. Beyond its theoretical relevance, the construction offers a pedagogically accessible framework for introducing core ideas in quantum mechanics and symmetry groups, making the abstract correspondence between $SU(2)$ and $SO(3)$ tangible through experimentally meaningful procedures.
Keywords: Quantum channels, Group homomorphism, Operational approach.
@@ I Introduction
Physical processes in quantum theory are described by completely positive and trace-preserving (CPTP) maps. By processes, we intend this term to be as broad as possible. Examples include measurement processes via quantum instruments [1], (ir)reversible time evolution [2], coarse-grained descriptions of blurred or saturated detectors [3], local operations and classical communication protocols composed of CPTP maps for applications such as entanglement theory [4], and the modeling of open quantum systems’ dynamics, where CPTP maps are employed to reproduce non-unitary evolution due to decoherence and possible dissipation [5]. This list can be extended indefinitely, reflecting the ubiquitous role of such maps in quantum theory and beyond — generalized probability theories, for example [6].
@@
Before the CP part, we start with the notion of only positivity, which is central to the mathematical and physical structure of quantum theory. In particular, positive operators are essential in the description of quantum state operators because positivity guarantees that all measurement probabilities are non-negative. Moreover, a density operator must be positive semidefinite to ensure that expectation values of all observables correspond to physically meaningful, real, and non-negative probabilities.
Going a little further, complete positivity is also a fundamental requirement for any physically admissible map. When a system interacts with an external environment, for instance, the joint evolution is described by a unitary transformation acting on the combined Hilbert space. After tracing out the environment, the reduced dynamics of the system are represented by a map acting solely on its state space. To ensure that this reduced map yields a valid quantum state not only for isolated systems but also when the system is part of a larger entangled one, the map must be completely positive. Without complete positivity, the extended action of the channel on entangled states could produce non-physical, non-positive density operators, violating the probabilistic consistency of quantum theory. A standard example is the transpose operation, which is positive but not completely positive [7]. For this reason, we also require the completely positive (CP) part.
Together, these two conditions—which define what is known as a quantum channel (or physical map)—appear to impose strong restrictions on the invertibility of CPTP maps with inverses that are still physically admissible. The depolarizing channel provides an example of a CPTP map whose inverse is not positive, highlighting the rarity — or even contingency — of finding a CPTP map with a CPTP inverse [8].
This question has already been studied in detail [8], where the characterization of invertible quantum channels is addressed via their Kraus representation. In essence, a physical map admits a CPTP inverse if and only if the Kraus decomposition of the former contains exactly one Kraus operator. Being a CPTP map, one demands that the Kraus representation is trace-preserving; this translates into the operators, say, $A_{a}$, with $a$ running in a finite set, adding up to identity, $\sum_{a}A^{*}_{a}A_{a}=\mathds{I}$. With just one operator in the representation, one concludes that the only operator is, actually, unitary.
In our contribution, we take this analysis a step further by exploring the connection between invertible CPTP maps with physical admissible inverse and group theory. Remarkably, the well-known two-to-one homomorphism between $SU(2)$ and $SO(3)$ can be recovered in a surprisingly simple way through a preparation procedure involving two-level quantum systems embedded in three-dimensional physical space.
@@
to emphasize that any two-level quantum system can, in practice, be prepared by selecting appropriate directions in three-dimensional physical space, as encoded operationally by the Bloch representation.
In this work we intentionally separate two distinct uses of the symbol $\mathds{R}^{3}$: (i) the physical space where laboratory procedures take place, and (ii) the Euclidean vector space that is used as a convenient representation space (for example the ambient space that contains the Bloch sphere). To avoid conflation we adopt an operational construction of physical space following the approach of [11, 12]: physical space is obtained from experimentally accessible procedures (marking points on rigid reference bodies, defining distances by compasses, forming classes under “firm unions” of rigid bodies, etc.), which yields a system of reference (SR) and an associated affine space $\mathcal{E}_{SR}$ together with its vector space of displacements $\mathds{D}$. This operational viewpoint emphasizes that points, displacements (vectors), directions and affine structure are defined by procedures and not assumed a priori.
With that operational scaffolding in place we can state the relation between Bloch vectors and the laboratory: Bloch vectors are elements of an abstract three-dimensional real vector space (the space in which the Bloch sphere $S^{2}$ is embedded), and, when a specific physical implementation is chosen, these vectors acquire an operational meaning as directions that label preparation procedures (for instance the orientation of a Stern–Gerlach magnetic field [13] or the settings of an $SU(2)$ polarization gadget [9]). The Hopf fibration provides the mathematical bridge: indistinguishable pure state rays in the normalized Hilbert sphere $S^{3}\subset\mathds{C}^{2}$ are projected onto points of $S^{2}\subset\mathds{R}^{3}$, so that each point on the Bloch sphere corresponds operationally to the family of laboratory preparations that produce that same state up to a global phase. The operational recipe for realizing the spherical coordinates $(\theta,\varphi)$ is therefore apparatus-dependent (e.g. SG magnet orientation or the Euler-angle settings of the universal $SU(2)$ polarization gadget), but the Hopf map guarantees a system-independent identification between the equivalence class of preparations and a Bloch direction.
Building on this, we go further to investigate transformations between vectors in $\mathds{D}$ (which give rise to quantum states in the sense explained above) and the corresponding allowed state transformations in the space of states. The sequence of these transformations is expected to make the diagram in Fig. 1 commute.
@@
On the one hand, an allowed transformation of a vector in physical space is represented by a rotation, which in turn induces a unitary transformation of the corresponding quantum state. This gives rise to the direction $SO(3)\rightarrow SU(2)$ in the homomorphism we aim to recover. It corresponds to the path represented by the black horizontal arrows in the diagram shown in Fig. 1. On the other hand, the reverse direction, $SU(2)\rightarrow SO(3)$, is obtained by requiring a reversible physical transformation of a quantum state. As previously discussed, the only permissible process of this kind is a unitary operation, which translates into a rotation in physical space. This step is illustrated by the blue arrows in the diagram.
To construct both directions in the diagram, we proceed as follows. In the next section, we show how a rotation in physical space unambiguously induces a unitary transformation on the quantum state associated with the corresponding vector. Section III provides a brief review of CPTP maps, with a focus on their Kraus representation. In particular, we discuss a characterization theorem for invertible CPTP maps with physically admissible inverses. Building on this result, in Section IV, we start from a unitary representing a reversible quantum process and reconstruct the associated rotation, thereby recovering the group two-to-one homomorphism we set out to rediscover. Finally, Section V is left for conclusions and some perspectives.
@@ II From physical spaces to state spaces: $SO(3)\rightarrow SU(2)$
Given a unit vector (called the Bloch vector) belonging to what is called the Bloch sphere $S^{2}$
$\vec{r}=(\sin\theta\cos\varphi,\sin\theta\sin\varphi,\cos\theta),$ (1)
with $\theta\in[0,\pi]$ and $\varphi\in[0,2\pi)$ in the physical space $\mathds{D}\cong\mathds{R}^{3}$, one constructs the pure state in $\mathds{C}^{2}$, or strictly speaking, in the projective space $\mathds{CP}^{1}$,
@@
The parameters $\theta$ and $\varphi$ may have a concrete physical interpretation. For example, in [13] it is shown that they indicate the direction of a magnetic field in the Stern-Gerlach experiment, preparing a bean of spin-up particles in that direction. Capitalizing in the Fig. 1, analogous to the isometry of $SU(2)$ and $SO(3)$, there is the isometry of the complex projective space $\mathds{CP}^{1}$ — the space of pure states of a qubit, modulo the global phase — with the Bloch sphere $S^{2}$ [14]. We restrict ourselves to the former.
@@
With this result in hand, we turn to transformations in physical space that preserve the nature of the quantum state—whether pure or mixed. Such transformations must be isometries, and can be described as rotations around a given axis $\hat{n}$ by an angle $\alpha$, denoted $\mathcal{R}(\hat{n},\alpha)$.
@@
Given a general rotation element $\mathcal{R}(\hat{n},\alpha)\in SO(3)$. The state $\rho_{\vec{r}^{\prime}}$, described by the vector $\vec{r}\,^{\prime}=\mathcal{R}(\hat{n},\alpha)\vec{r}$, can be calculated by applying a matrix $\mathcal{U}(\hat{n},\alpha)$ $\in SU(2)$ on the state $\rho_{\vec{r}}$, described by the original vector $\vec{r}$, in such a way that $\rho_{\vec{r}^{\prime}}=\mathcal{U}(\hat{n},\alpha)\rho_{\vec{r}}\,\mathcal{U}^{*}(\hat{n},\alpha)$.
@@
$\mathcal{R}_{jk}(\hat{n},\alpha)$ $=\delta_{jk}\cos\alpha+n_{j}n_{k}(1-\cos\alpha)-\sum^{3}_{l=1}\varepsilon_{jkl}n_{l}\sin\alpha$
@@
and the group generators $\vec{\tau}$ are given by $(\tau_{l})_{jk}=-i\varepsilon_{jkl}$ [17]. From now on, Latin letters ($i,j,k,$...), as customary, assume the values $1,2,3$.
Now, if $\vec{r}\,^{\prime}=\mathcal{R}(\hat{n},\alpha)\vec{r}$, then the rotated vector $\vec{r}\,^{\prime}$ gives rise to a state $\rho_{\vec{r}^{\prime}}$ that we seek to determine. We have
$\rho_{\vec{r}^{\prime}}$ $=\frac{1}{2}\left(\mathds{I}+\vec{r}\,^{\prime}\cdot\vec{\sigma}\right)=\frac{1}{2}\left(\mathds{I}+(\mathcal{R}\vec{r})\cdot\vec{\sigma}\right)$
$=\frac{1}{2}[\mathds{I}+\cos\alpha(\vec{r}\cdot\vec{\sigma})+(1-\cos\alpha)(\hat{n}\cdot\vec{r})(\hat{n}\cdot\vec{\sigma})$
$-\sin\alpha(\vec{r}\times\hat{n})\cdot\vec{\sigma}].$ (5)
In the expression above, we have used the explicit action of (4) on the vector $\vec{r}$, that is,
$\vec{r}\,^{\prime}=\cos\alpha\vec{r}+(1-\cos\alpha)(\hat{n}\cdot\vec{r})\hat{n}-\sin\alpha\,\vec{r}\times\hat{n}$. (6)
Using the trigonometrical identities
$\cos\alpha=\cos^{2}{\frac{\alpha}{2}}-\sin^{2}{\frac{\alpha}{2}},\quad\,\,\sin\alpha=2\sin{\frac{\alpha}{2}}\cos\frac{\alpha}{2},$
and
$1-\cos\alpha=2\sin^{2}{\frac{\alpha}{2}}$
one finds that
$\vec{r}\,^{\prime}\cdot\vec{\sigma}=\left(\cos^{2}{\frac{\alpha}{2}}-\sin^{2}{\frac{\alpha}{2}}\right)(\vec{r}\cdot\vec{\sigma})$-
$-2\sin{\frac{\alpha}{2}}\cos{\frac{\alpha}{2}}(\vec{r}\times\hat{n})\cdot\vec{\sigma}+2\sin^{2}{\frac{\alpha}{2}}(\hat{n}\cdot\vec{r})(\hat{n}\cdot\vec{\sigma}).$
Rearranging terms, we have
$\vec{r}\,^{\prime}\cdot\vec{\sigma}$ $=\cos^{2}{\frac{\alpha}{2}}(\vec{r}\cdot\vec{\sigma})-2\sin{\frac{\alpha}{2}}\cos{\frac{\alpha}{2}}(\vec{r}\times\hat{n})\cdot\vec{\sigma}$
$+\sin^{2}{\frac{\alpha}{2}}(2(\hat{n}\cdot\vec{r})(\vec{r}\cdot\vec{\sigma})-(\vec{r}\cdot\vec{\sigma})).$
From $(\vec{r}\cdot\vec{\sigma})(\vec{n}\cdot\vec{\sigma})=(\vec{r}\cdot\hat{n})\mathds{I}+i(\vec{r}\times\hat{n})\cdot\vec{\sigma}$, the Pauli product identity, we can extract two identities, namely,
$(\hat{n}\cdot\vec{\sigma})(\vec{r}\cdot\vec{\sigma})(\hat{n}\cdot\vec{\sigma})=2(\hat{n}\cdot\vec{r})(\vec{r}\cdot\vec{\sigma})-(\vec{r}\cdot\vec{\sigma});$
$(\vec{r}\cdot\vec{\sigma})(\hat{n}\cdot\vec{\sigma})-(\hat{n}\cdot\vec{\sigma})(\vec{r}\cdot\vec{\sigma})=2i(\vec{r}\times\hat{n})\cdot\vec{\sigma}.$
Knowing this, our $\vec{r}\,^{\prime}\cdot\vec{\sigma}$ term than becomes
$\vec{r}\,^{\prime}\cdot\vec{\sigma}$ $=\cos^{2}{\frac{\alpha}{2}}(\vec{r}\cdot\vec{\sigma})-i\cos{\frac{\alpha}{2}}\sin{\frac{\alpha}{2}}(\hat{n}\cdot\vec{\sigma})(\vec{r}\cdot\vec{\sigma})+$
$+i\cos{\frac{\alpha}{2}}\sin{\frac{\alpha}{2}}(\vec{r}\cdot\vec{\sigma})(\hat{n}\cdot\vec{\sigma})+$
$+\sin^{2}{\frac{\alpha}{2}}(\hat{n}\cdot\vec{\sigma})(\vec{r}\cdot\vec{\sigma})(\hat{n}\cdot\vec{\sigma}),$
and finally, with $\mathcal{U}=\cos{\frac{\alpha}{2}}\mathds{I}-i(\hat{n}\cdot\vec{\sigma})\sin{\frac{\alpha}{2}}$, we can rearrange the expression above to
$\vec{r}\,^{\prime}\cdot\vec{\sigma}=\mathcal{U}\ (\vec{r}\cdot\vec{\sigma})\ \mathcal{U}^{*}.$ (7)
The expression (7) allows us to rewrite our $\rho_{\vec{r}^{\prime}}$ from equation (5) as
$\rho_{\vec{r}^{\prime}}=\mathcal{U}(\hat{n},\alpha)\rho_{\vec{r}}\,\mathcal{U}^{*}(\hat{n},\alpha)$ (8)
@@
$\mathcal{U}(\hat{n},\alpha)=\cos\frac{\alpha}{2}\mathds{I}-i\sin\frac{\alpha}{2}(\hat{n}\cdot\vec{\sigma})=e^{-i\frac{\alpha}{2}\hat{n}\cdot\vec{\sigma}},$ (9)
@@
The result above shows that each rotation $\vec{r}\,^{\prime}=\mathcal{R}(\hat{n},\alpha)\vec{r}$, with $\mathcal{R}(\hat{n},\alpha)=e^{-i\alpha\hat{n}\cdot\vec{\tau}}$ induces a state transformation $\rho_{\vec{r}^{\prime}}=\mathcal{U}(\hat{n},\alpha)\rho_{\vec{r}}\,\mathcal{U}^{*}(\hat{n},\alpha)$ with $\mathcal{U}(\hat{n},\alpha)=e^{-i\frac{\alpha}{2}\hat{n}\cdot\vec{\sigma}}$. Hence we have just established the mapping $\phi:SO(3)\rightarrow SU(2)$, with $\phi(\mathcal{R}(\hat{n},\alpha))=\mathcal{U}(\hat{n},\alpha)$. It is important to emphasize a peculiarity of this mapping. Since in (8) the unitary operator $\mathcal{U}$ and its conjugate appear jointly, the elements $\mathcal{U}$ and $\mathcal{U}$ yield the same physical result. Therefore, (9) does not, strictly speaking, define a mapping from spatial rotations to elements of $SU(2)$, but rather from the set of parameters $(\vec{n},\alpha)$ to $SU(2)$. Additional remarks concerning this behavior will be provided in Section IV, specifying the double-cover structure of $SU(2)$ over $SO(3)$.
Our next step will be focused on constructing the inverse of $\phi$. To that end, we will make a short digression on CPTP maps.